A Package to Put Figures and Tables at the End of an Article

Patrick W. Daly

This paper describes package figcaps version 4.7 from 1999/02/23 It is part of the preprint collection of packages

Summary

The stripped version of this file contains the following brief description:

%  This option allows the figure captions to be collected throughout
%    the paper and printed on a separate page at the end. The figures
%    themselves will not appear in the text. This is for purposes of
%    a manuscript for submission.
%  Similarly, tables are not printed in the text, but are outputted at
%    the end, after the figure captions.
<*plates>
%  A plate environment also exists, handled just like figures.
</plates>
<*longtab>
% A longtable environment exists, for use with the longtable tools package.
</longtab>
%  The figures themselves may appear, with captions, at
%    the end, after the tables. This is done with command \printfigures
%    in the preamble. This only makes sense if the figures have been
%    generated by LaTeX or if they are being imported.
%  Marginal notes added where environments placed if \figmarkon given.
%  NOTE: Only works for article-like styles at present, where \section is
%        the highest sectioning level.
<*!toc>
%  NOTE: Tables of contents (incl figures and tables) cannot be used
</!toc>
%  NOTE: This file must be read in AFTER \appendix has been defined;
<*sublab>
%        if sublabel.sty is also used, it must be read in first.
</sublab>
<agu> % The captions are formatted as for AGU journals.
<egs> % The captions are formatted as for EGS journals.
<*!209>
%
% Options with LaTeX2e:
%   printfigures  - prints the figures at the end of the article, (is
%                   like issuing the \printfigures command)

 %   figmark       - turns on the markers for figures and tables in the text,
 %                   (is like issuing \figmarkon)
 %   figon/off     - activates or deactivates the whole procedure (is like
 %                   (issuing \figcapson or \figcapsoff). Default=on.
 %   mylang        - (default) leaves explicit text (e.g. `Figure Captions'
 %                   as is, or if undefined, makes them English
 %   english,american,german,austrian,french,esperanto - changes explicit
 %                   texts accordingly
 %   blank         - makes explicit texts empty.
</!209>

1 Introduction

Often when producing manuscripts for submission to a journal, it is required that the figure captions be listed at the end on a separate page, and also that tables be put at the end too. Then when the manuscript is ready and camera-ready copy, or a preprint, is to be generated, one has to move them all back to the proper positions in the paper.

This style alleviates this problem. The figure and table environments are placed in the text as they should be for the preprint or camera-ready copy, but the tables actually are printed at the end. The figures are not printed at all, but their captions are listed on a separate page at the end. Optionally, the figures themselves may be printed at the very end. This last only makes sense if there is something in the figures, i.e., they have been produced with picture or are imported graphics.

2 Invoking the Package

The macros in this package are included in the main document with the \usepackage command of L^{A}TEX 2ε,

     \documentclass[..]{...}
     \usepackage[options]{figcaps}

where the possible options are described below.

Alternatively, the name of the package is added as an option to the \documentstyle command in L^{A}TEX 2.09 compatibility mode, as

     \documentstyle[..figcaps..]{...}

It may only be used for article, or an article-like main class. That is, for one that does not contain the \chapter level of sectioning.

The order in which other packages are loaded is important for certain features. If my package sublabel is included, then it must come before figcaps. Furthermore, figcaps adds to the existing definition of \appendix, so that if the author has any packages that alter this definition, they must come before figcaps. Clearly, \appendix may not be redefined afterwards.

3 Using the Package

\figcapson The macro package in figcaps operates pretty much automatically. The whole \figcapsoff thing may be turned on and off with the commands \figcapson and \figcapsoff.

The default is ⟨on⟩. One might want to control the placement of figures and tables with this flag rather than by adding and removing the option from the

\documentclass command.

Alternatively, the L^{A}TEX 2ε options figon and figoff have the same effect, with figon being the default.

\printfigures The only other means of interacting with the macros is with the command \printfigures, which, when given anywhere in the text, enables the figures themselves to be output at the very end of the paper. Without this, only the captions are listed. With it, both captions and figures are output.

The L^{A}TEX 2ε option printfigures in \usepackage is equivalent to issuing the command \printfigures in the text.

\figmarkon Another option is to put marginal notes into the text at the place where the \figmarkoff float environment is entered. This is useful for camera-ready copy that will be cut-

and-pasted together to indicate where the figures and tables should appear. The commands \figmarkon and \figmarkoff control this feature, with the default being ⟨off ⟩.

The L^{A}TEX 2ε option figmark in \usepackage sets this default to ⟨on⟩; in other words, it is equivalent to issuing \figmarkon in the text.

\figurecapname The text that is written as header at the top of the special pages is defined

\tablepagename
\figurepagename

in three commands \figurecapname, \tablepagename, and \figurepagename.

They default to English text, but may be redefined to other languages, or suppressed altogether, with \renewcommand or with L^{A}TEX 2ε options (see page 4).

4 How it Works

The contents of all the figure and table environments are written to auxiliary files .lof and .lot respectively, and then these are read in again at the end of the paper. Since these auxiliary files are normally used to write information for the lists of figures and tables, this capability is lost. However, it is not likely that one wants such lists, or even a table of contents, in a manuscript. In fact, the

commands \tableofcontents, \listoffigures, \listoftables are turned off

(they issue a warning message).

If you really do need tables of contents, then see the extra features below.

4.1 With the longtable Package

This package may optionally be used with the longtable environment defined in the package of the same name. Since such tables are normally long, i.e. over a page, the list of tables at the back will always start a new page for each of these.

For preprints, when the tables are not moved to the back (\figcapsoff), it makes sense to put the longtable environment inside the \afterpage command (from the afterpage package). However, with \figcapson, and the table is move, this is useless. In this case, the \afterpage command is deactivated. One should realize this in case it is used for other purposes.

4.2 As Module to Journal-Specific Styles

Although figcaps is meant to be an all-purpose bibliographic style package, it may also be incorporated as a module to other packages for specific journals. In this case, many of the general features may be left off. This is allowed for with docstrip options that not only leave off certain codelines, but also include extra ones. So far, options exist for

egs for European Geophysical Society journals.

agu for American Geophysical Union journals.

Previous options jgr and grl have become obsolete due to revisions in these journals; they have been replaced by the more general agu option.

5 L^{A}TEX 2ε Options

Under L^{A}TEX 2ε, options may be added to the \usepackage command that loads the package. Available options for figcaps set certain control flags, thus replacing the equivalent commands, or select the language for the texts that are printed by

\figurecapname, \tablepagename, and \figurepagename.

figon (default) activates the whole procedure of moving figures and tables to the back of the article, (same as \figcapson);

figoff deactivates the whole procedure, (same as \figcapsoff);

printfigures allows the figures themselves and not just their captions to be output at the very end, in addition to the caption page, (same as \printfigures); by default this is not done;

figmark enables marking of figure and table environments in the text with marginal notes; (same as \figmarkon);

mylang (default) leaves the three name commands as they are; however if they are not defined at all, then they are given English values; since these name commands are not standard, one cannot expect them to be set by other language adaption packages;

english, american, german, austrian, french, esperanto set the name com-

mands to the translation in the corresponding language, overwriting any previous definitions;

blank sets the text of \figurecapname, \tablepagename, and \figurepagename

to nothing.